An escalator is used to move 15 passengers every minute from the first floor of a department store to the second. The second floor is located 6 meters above the first floor. The average passenger's mass is 52 kg. Determine the power requirement of the escalator in order to move this number of passengers in this amount of time. *


1
Expert's answer
2022-02-23T08:30:31-0500

The power is given by

"P=\frac{W}{t}=\frac{Nmgh}{t}=\rm \frac{15*52*9.8*6}{60}=764\: W"